OpenStudy (anonymous):

Tickets to Friday’s football game are $6.00 for adults and $2.00 for students. If the school pre-sold 58 tickets for $216, how many adult and student tickets were purchased? 30 adults, 28 students 25 adults, 33 students 42 adults, 16 students 20 adults, 38 students @iGreen

OpenStudy (anonymous):

B :P

OpenStudy (anonymous):

your too slow igreen :D

OpenStudy (igreen):

This is just like the last question. You can use these two equations: 6x + 2y = 216 x + y = 58 Where x = Adults And y = Students Solve and you get (25, 33) So that's 25 Adults and 33 Students.

OpenStudy (igreen):

@J-bird Because you're giving out direct answers without explanation WHICH YOU'RE NOT SUPPOSED TO DO. If I did what you did I'd have this question answered in 2 seconds.

OpenStudy (igreen):

@ganeshie8 @Preetha This is the third warning I gave you.

OpenStudy (anonymous):

what do you think i did??

OpenStudy (anonymous):

what grade are you in igreen??

OpenStudy (anonymous):

Tickets to Friday’s football game are $5.50 for adults and $4.50 for students. If the school pre-sold 25 tickets for $124.50, how many adult and student tickets were purchased? 7 adults, 18 students 10 adults, 15 students 12 adults, 13 students 15 adults, 10 students

OpenStudy (anonymous):

hes gone, but i help you on this one too

OpenStudy (anonymous):

k

OpenStudy (anonymous):

this is a different one ^^ @J-bird

OpenStudy (igreen):

I'm in 10th grade.

OpenStudy (anonymous):

12 adults and 13 students

OpenStudy (anonymous):

im 9th

OpenStudy (anonymous):

and there not a direct answer this time :P

OpenStudy (anonymous):

Solve by elimination. 2x − y = 5 3x − 2y = 6 (0, negative-3) (3, 4) (4, 3) (1, 4)

OpenStudy (igreen):

Yes it is. How is "12 adults and 13 students" not a direct answer?

OpenStudy (anonymous):

i didnt say C

OpenStudy (igreen):

http://openstudy.com/code-of-conduct Give Help, Not Answers I will encourage and guide those needing help, and not just give them an answer OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. Have homework problems? We can help you - as long as you are looking to learn how to get your solution (and not just the answer).

OpenStudy (igreen):

Yeah, you just copy and pasted one of the options.

OpenStudy (anonymous):

Solve by elimination. 2x − y = 5 3x − 2y = 6 (0, negative-3) (3, 4) (4, 3) (1, 4)

OpenStudy (anonymous):

you think i copy and pasted the answers??

OpenStudy (igreen):

You have to EXPLAIN WHY YOU GOT THE ANSWER You don't just give the answer.

OpenStudy (igreen):

If you DON'T EXPLAIN YOUR ANSWER, then you're giving a DIRECT ANSWER.

OpenStudy (anonymous):

if they want an explanation i would give it to them

OpenStudy (igreen):

Obviously you must be in 2nd grade, because you don't know what a "direct answer" is. A direct answer is not saying A, B, C, or D only, it's giving them the answer without explaining.

OpenStudy (igreen):

No, it's not their choice whether or not you give them an explanation, you HAVE TO give them an explanation. Otherwise they will never know how to do it in the future.

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

chill, just chill :P